gababb09:

The mass of a freight train is 9.98 ⋅ 10^6 kilograms. The mass of an aircraft carrier is 7.3 ⋅ 10^7 kilograms. Approximately how many more kilograms is the mass of an aircraft carrier than the mass of a freight train? Show your work and write your answer in scientific notation.

Hoodmemes:

Well u wanna do the first part for the freight train \[9.98\times10^6 =9,980,000\] Second part for the aircraft carrier \[7.3\times10^7=73,000,000\] Now u wanna subtract the freight train by the aircraft carrier \[73,000,000- 9,980,000=63,020,000\] Concluding in the scientific notation being: \[6.302 × 10^7\] There are approximately 6.302 x 10^7 many more kilograms is the mass of an aircraft carrier than the mass of a freight train when put in the form of scientific notation

Hoodmemes:

Does this help?

gababb09:

yes it does thanks Harry much I have 2 more questions that I will post in a few minutes do you think you can help answer them they are simular to this one

Hoodmemes:

oh ok